Single Plantwide Factory Overhead Rate

Spotted Cow Dairy Company manufactures three products—whole milk, skim milk, and cream—in two production departments, Blending and Packing. The factory overhead for Spotted Cow Dairy is $352,800.

The three products consume both machine hours and direct labor hours in the two production departments as follows:

Direct Labor Hours Machine Hours
Blending Department
Whole milk 250 630
Skim milk 270 580
Cream 220 230
740 1,440
Packing Department
Whole milk 310 360
Skim milk 490 450
Cream 140 150
940 960
Total 1,680 2,400

Required:

1. Determine the single plantwide factory overhead rate, using each of the following allocation bases: (a) direct labor hours and (b) machine hours. If required, round all amounts to the nearest dollar.

a. Direct labor overhead rate $ per direct labor hour
b. Machine hour overhead rate $ per machine hour

2. Determine the product factory overhead costs, using (a) the direct labor hour plantwide factory overhead rate and (b) the machine hour plantwide factory overhead rate.

Whole Milk Skim Milk Cream
Direct labor hours $ $ $
Machine hours $ $ $
Answer #1

Single Plantwide Rate = Total Overheads/Total labor hours

= 352800/1680

= $210 per direct labor hour

Machine hour overhead rate = 352800/2400

= $147 per machine hour

Overhead = Rate*Driver consumed

Whole Milk

Skim Milk

Cream

Direct labor hours

       117,600

       159,600

       75,600

Machine hours

       145,530

       151,410

       55,860

В. А. 1 Overhead = Rate* Driver consumed C D 2 3 Direct labor hours 4 Machine hours Whole Milk Skim Milk Cream =(250+310)*210